You are here: Foswiki>PartnerObserving Web>TechNightLogs>September24UT (21 Jul 2009, NormCushing)Edit Attach
Afternoon

  • State of things: Spare controller which was swapped in earlier last night remains in place (ID #1). But we got the same error message about a bad bias voltage with the spare as with the original controller. Fernando & Vincenzo noted a bias cable with a bad pin - they replaced this but it was not the cause of the problem. Also they note that the saturation level is <65000 ADU for some chips.

  • The problem we had with the bad bias voltage seems to be from the controller software and appears when an image is taken after one which uses only a single chip. So any image taken after a 'fastextra' which uses only chip 2, e.g., would fail with the error. Apparently the bias voltage is not reset correctly after the single chip image is taken.

  • Andrew examined a set of superfoc images taken last night (the one starting with lbcb.20080923.120955) and noted a focal plane tilt around the E-W axis (for PA=0). I had noticed this the night before in a superfoc sequence and tried to obtain a sequence at PA=180 as well as PA=0, but, because of a focus offset, probably due to poor collimation, results were not conclusive. We need to try to obtain PA=0 and PA=180 focus sequences to confirm.

  • Last night's test with adjusting Z11 to remove the ring helped lbcfpia converge to the proper focus. After lbcfpia converged, the resulting pupil image still had a bright ring. Z11 = -400 was manually added to remove the ring, and Z4 = +500 added to correct the pupil size, so that the outer diameter would have the nominal value of ~ 60.5pix. Resulting in-focus images had fwhm ~ 2.5 pixels at V. Note there was still on-axis astigmatism apparent. In fact lbcfpia results indicated ~200-300nm astigmatism. We need to try to bring this number down to ~100nm.

Evening

  • 19:30MST Going to replace resistor on blue window heater. There was some condensation on the blue dewar window again tonight.

  • Fernando & Vincenzo noted that Chip2 and Chip4 of LBC-Red saturate before 65000 - Chip2 at 12000 ADU and Chip4 at 20000 ADU. This is not a hard saturation but rather appears as a loss of sensitivity. They are working to fix this.

  • 20:40MST Vincenzo modifying OBs for fastextra to use all 4 chips,but windowed. There is no modification needed for lbcfpia.pro, however the observer must add some arguments to the command line. For blue, run lbcfpia with the syntax: lbcfpia, LBCChip=2, ImgSec=[2403,4450,3408,4608]

  • 21 Turn on camera with telescope at zenith and red fails with bad bias voltage - so the hypothesis that it failed only when a four-chip image was taken after a single-chip image is sunk. [later ... no -this was understood].

  • 23:15 David is on the phone with Paul G. since there is a problem with the telescope. The servers are intermittently hanging(?). LBC observers have been testing the red channel.
  • 23:30 David is stopping and restarting TCS.
  • 23:45 Fernando talking to Andrea diPaola. Andrea will try to patch sw to make single chip not leave it in a bad state. TCS problem is fixed. Trying to connect LBC to telescope and will observe with blue/red with caveat that red saturation levels are wrong for chips 2 & 4.
  • 00:15 Telescope is back up and running.
  • 01:00 PartnerObserving with blue and red, but red still has the 'saturation' problem - i.e. stars saturate on chip2 at 12000 ADU and on chip4 at 25000 ADU instead of the normal 65000 ADU. Image shows fwhm ~ 4.7pix (~1.1") in Blue. Seeing does not appear as good as last night but we think this is real, not an artifact of bad focus, because pupils are fuzzy. Also, the Polaris monitor was showing ~1 last night but tonight is up around 4-5" - a bit of a steep climb, really.
  • 04:00 Started a new OB but got a warning on the Red camera - rotator failure. Replayed but got a similar warning on the Blue camera - rotator. Loaded and started another OB - again a rotator error. In the log there is also a message about Red camera controller error code 2. Turned off/on LBC - it came up fine. Note that elevation was ~45deg. This is issue #1664.
  • 04:45 Red controller error at the end of the OB. Log shows 'error code 2', as before. But also there are many temperature warnings. All red systems 'stop'ed. Vincenzo stopped/restarted camera but red will not come up because of problems reading the temperature. 2008/09/24 11:49:21.024763 W R HKEEPING TEMPERAT could not read temperature value [src/housekeeping/housekeeping.c:1138] This is issue #1663.

  • 04:49 Polaris monitor showing ridiculous seeing - seems to be increasing and is ~6-7". Images are about 1.2"
  • 05:40 around this time took sky flats
  • 05:51 pretty light out - closing

Things to be aware of
  • IDL prompt does not return after lbcfpia sends corrections to telescope. The mirror moves as expected, but the lbcfpia or TCSSendWavefront program hangs with the message about sending to TCS...
  • z-SLOAN filter on LBC-Red has some debris on it. Vincenzo has pictures of it. Work right now has been focussed on red controller, so this has not been investigated further.
  • Blue camera: Background on chips 3 and 4 shows a scalloped pattern. This is constant and comes out in the flat field. It resulted from a solvent or cleaner used during the work on the chips last summer.
  • For Blue camera, a manual correction of Z8 = 2000nm has been needed at the start of the night.
  • Focus procedure to remove 'ring' around pupil outer edges and help LBCFPIA achieve accurate results:
    • Focus/Collimate with LBCFPIA
    • Collimate with LBCFPIA
    • If the resulting 'collimated' pupils shows a ring, enter Z11 to remove the ring. Z11 = -400nm was used last night.
    • Adjust Z4 so that the outer diameter is 60.5 pixels. You can measure this with 'regions, ruler' command on ds9.
    • Subsequent runs of LBCFPIA should not show ring and will result in better focus. Focus value when there is a ring present may be off by 50-100 microns.
  • There is a new GUI version - it is in /home/lbcobs/GUIv2.0_20080923/ .
-- OlgaKuhn - 23 Sep 2008
Topic revision: r8 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback